OpenStudy (anonymous):

GIVNING MEDALS Jenny is putting money into a savings account. She starts with $350 in the savings account, and each week she adds $60. Let S represent the total amount of money in the savings account (in dollars), and let W represent the number of weeks Jenny has been adding money. Write an equation relating S to W. Then use this equation to find the total amount of money in the savings account after 11 weeks. Equation: ??? Total amount of money after 11 weeks??

OpenStudy (anonymous):

@ganeshie8 @phi @katkipe @lovelychap @ogthot @edded please help

OpenStudy (anonymous):

@goalieboy @party_girl @Tazmaniadevil @bohotness @jakewoln @josedavid @YanaSidlinskiy @Rainbow_dashie @KlOwNlOvE @goformit100

OpenStudy (josedavid):

what grade is this

OpenStudy (anonymous):

12TH

OpenStudy (josedavid):

cool

OpenStudy (anonymous):

I just need an equation , and i can solve the rest

OpenStudy (anonymous):

Okay I think the equation would be 350s + 60w = x (x will be the total)

OpenStudy (anonymous):

Let me test it, I may be wrong.

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

oh wait no 350 + 60w = s

OpenStudy (anonymous):

you're putting 60 in per week and w represents the number of weeks those will be multiplied s is the total cash if 350 is already in there it will be added into the equation

OpenStudy (phi):

****I just need an equation , and i can solve the rest **** The whole point of this question is to be able to figure out the equation. If you can't, then you should work at learning how to.

OpenStudy (anonymous):

phi is right lol T_T

OpenStudy (anonymous):

@KlOwNlOvE how would i do that

OpenStudy (anonymous):

All of the operations are in my statement you don't know what S is yet because you haven't solved for 11 weeks so S will be on one side all the other numbers and variables will be on the other

OpenStudy (anonymous):

so is it like this --> 60w+350=s ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay how do i solve it now , which is first

OpenStudy (anonymous):

I was right c:

OpenStudy (anonymous):

substitute 11 in for S then follow your order of operations starting with multiplying then finish with adding S=?

OpenStudy (anonymous):

sorry 11 in for w

OpenStudy (anonymous):

so 60(11)+350=s ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is it 1010?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thankk you @KlOwNlOvE

OpenStudy (anonymous):

you're welcome